Consider the linear eigenproblem Ax=λx for the matrix

D=[1 1 2

2 1 1

1 1 3 ]

1. Solve for the largest (in magnitude) eigenvalue of the matrix and the corresponding eigenvector by the power method with 𝑥(0)T=[1 0 0]

2. Solve for the smallest eigenvalue of the matrix and the corresponding eigenvector by the inverse power method using the matrix inverse. Use Gauss-Jordan elimination to find the matrix inverse.


1
Expert's answer
2021-10-26T03:18:59-0400

1.

"x_1=Ax_0=\begin{pmatrix}
 1 & 1&2 \\
 2 & 1&1\\
1&1&3
\end{pmatrix}\begin{pmatrix}
 1 \\
 0 \\
0
\end{pmatrix}=\begin{pmatrix}
 1 \\
 2 \\
1
\end{pmatrix}"


"x_2=Ax_1=\begin{pmatrix}
 1 & 1&2 \\
 2 & 1&1\\
1&1&3
\end{pmatrix}\begin{pmatrix}
 1 \\
 2 \\
1
\end{pmatrix}=\begin{pmatrix}
 5 \\
 5 \\
6
\end{pmatrix}"


"x_3=Ax_2=\begin{pmatrix}
 1 & 1&2 \\
 2 & 1&1\\
1&1&3
\end{pmatrix}\begin{pmatrix}
 5 \\
 5 \\
6
\end{pmatrix}=\begin{pmatrix}
 22 \\
 21 \\
28
\end{pmatrix}"


"x_4=Ax_3=\begin{pmatrix}
 1 & 1&2 \\
 2 & 1&1\\
1&1&3
\end{pmatrix}\begin{pmatrix}
 22 \\
 21 \\
28
\end{pmatrix}=\begin{pmatrix}
 99 \\
 93\\
127
\end{pmatrix}"


"x_5=Ax_4=\begin{pmatrix}
 1 & 1&2 \\
 2 & 1&1\\
1&1&3
\end{pmatrix}\begin{pmatrix}
 99 \\
 93 \\
127
\end{pmatrix}=\begin{pmatrix}
 446 \\
 418\\
573
\end{pmatrix}=573\begin{pmatrix}
 0.78 \\
 0.73\\
1
\end{pmatrix}"


dominant eigenvector:

"x=\begin{pmatrix}
 0.78 \\
 0.73\\
1
\end{pmatrix}"


"Ax=\begin{pmatrix}
 1 & 1&2 \\
 2 & 1&1\\
1&1&3
\end{pmatrix}\begin{pmatrix}
 0.78 \\
 0.73\\
1
\end{pmatrix}=\begin{pmatrix}
 2.51 \\
 4.75\\
10.26
\end{pmatrix}"


"Ax\cdot x=15.69"

"x\cdot x=2.14"


dominant eigenvalue:

"\lambda=\frac{Ax\cdot x}{x\cdot x}=\frac{15.69}{2.14}=7.33"


2.

"\begin{pmatrix}
 1 & 1&2 &|1&0&0\\
 2 & 1&1&|0&1&0\\
1&1&3&|0&0&1
\end{pmatrix}\to\begin{pmatrix}
 1 & 0&0 &|-2&1&1\\
 0 & 1&0&|5&-1&-3\\
0&0&1&|-1&0&1
\end{pmatrix}"


"A^{-1}=\begin{pmatrix}
 -2& 1&1 \\
 5 & -1&-3\\
1&0&1
\end{pmatrix}"


"x_1=A^{-1}x_0=\begin{pmatrix}
 -2& 1&1 \\
 5 & -1&-3\\
1&0&1
\end{pmatrix}\begin{pmatrix}
 1 \\
 0 \\
0
\end{pmatrix}=\begin{pmatrix}
 -2 \\
 5 \\
1
\end{pmatrix}"


"x_2=A^{-1}x_1=\begin{pmatrix}
 -2& 1&1 \\
 5 & -1&-3\\
1&0&1
\end{pmatrix}\begin{pmatrix}
 -2 \\
 5 \\
1
\end{pmatrix}=\begin{pmatrix}
 2 \\
 2 \\
-1
\end{pmatrix}"


"x_3=A^{-1}x_2=\begin{pmatrix}
 -2& 1&1 \\
 5 & -1&-3\\
1&0&1
\end{pmatrix}\begin{pmatrix}
 2 \\
 2 \\
-1
\end{pmatrix}=\begin{pmatrix}
 -3 \\
 11 \\
1
\end{pmatrix}"


"x_4=A^{-1}x_3=\begin{pmatrix}
 -2& 1&1 \\
 5 & -1&-3\\
1&0&1
\end{pmatrix}\begin{pmatrix}
 -3 \\
 11 \\
1
\end{pmatrix}=\begin{pmatrix}
 18 \\
 -29 \\
-2
\end{pmatrix}"


"x_5=A^{-1}x_4=\begin{pmatrix}
 -2& 1&1 \\
 5 & -1&-3\\
1&0&1
\end{pmatrix}\begin{pmatrix}
 18 \\
 -29 \\
-2
\end{pmatrix}=\begin{pmatrix}
 -67 \\
 125 \\
16
\end{pmatrix}=16\begin{pmatrix}
 -4.19 \\
 7.81 \\
1
\end{pmatrix}"


the corresponding eigenvector the smallest eigenvalue:

"x=\begin{pmatrix}
 -4.19 \\
 7.81 \\
1
\end{pmatrix}"


"A^{-1}x=\begin{pmatrix}
 -2& 1&1 \\
 5 & -1&-3\\
1&0&1
\end{pmatrix}\begin{pmatrix}
 -4.19 \\
 7.81 \\
1
\end{pmatrix}=\begin{pmatrix}
 17.19 \\
 -31.76 \\
-3.19
\end{pmatrix}"


the smallest eigenvalue:

"\lambda=\frac{x\cdot x}{A^{-1}x\cdot x}=\frac{26.37}{-323.26}=-0.08"